# High thermal conductivity dominated by thermal phonons with mean free paths exceeding hundred nanometer in three-dimensional covalent organic framework derivatives: a molecular dynamics study

### v1

@@ I Introduction
Covalent organic frameworks (COFs), a class of crystalline porous organic materials, hold promise for various applications such as gas storage and separation Furukawa and Yaghi (2009); Fan et al. (2018), catalysis Wang et al. (2018), optoelectronics Keller and Bein (2021) and energy storage Zhu et al. (2021), owing to their large surface areas, tunable pore sizes, and high thermal stabilities Côté et al. (2005); El-Kaderi et al. (2007); Diercks and Yaghi (2017). Prior extensive studies have suggested that the diverse nature of structure, composition, and functionalities facilitates the tunability of their physical properties for tailored applications Huang et al. (2016); Liu et al. (2021). For instance, conjugated COFs have shown unique optoelectronic phenomena such as exceptionally high charge carrier mobility Guo et al. (2013) and tunable fluorescence emission as a result of cation addition Ding et al. (2016). Additionally, COFs could exhibit other interesting properties such as facile ion transport Wang et al. (2017); Li et al. (2021) and thermoelectricity Wu et al. (2023); Zhou et al. (2023), which are crucial for energy storage and conversion applications. However, efficient heat dissipation is needed in such applications to ensure device stability, making it essential to understand heat transport properties in COFs Shakouri (2006); Pop (2010); Babaei et al. (2020).
@@
II Results and discussion
We studied the thermal transport of COF-102 derivatives using MD simulations. Detailed descriptions of the MD simulations implemented in this work are given in Supporting Information Sec. SI and III Methods. As illustrated in Fig. 1(a), the COF-102 derivatives studied in this work were constructed from two cubic topologies, carbon nitride (ctn) and boracite (bor). They are formed by linking tetrahedral carbon nodes with 3-connected linkers; these nodes are connected to three distinct linkers via C-C bonds. The size of linker depends on the number of phenyl rings it contains. Although both ctn and bor topologies share the same stoichiometry, bor topologies are generally % less dense than ctn topologies (see Supporting Information Sec. SII for details).
Figure 1(b) shows room temperature thermal conductivities of COF-102 derivatives, calculated using the Green-Kubo formalism. As expected, a positive correlation between thermal conductivity and mass density is clearly visible. We note that the opposite trend is observed between thermal conductivity and pore size (see Supporting Information Fig. S3 in Sec. SI). COF-102-1, in particular, exhibits the highest thermal conductivity ( Wm-1K-1; Wm-1K-1), while the others show thermal conductivities reduced by nearly two orders of magnitude. In all cases, bor topologies yield % lower thermal conductivities than ctn topologies. We here note that COF-102-2 (ctn) is structurally similar to the original COF-102, except that its linkers are formed by substituting all atoms in the boroxine (B3O3) rings with carbon atoms. This leads to enhanced thermal conductivity of Wm-1K-1, which is % larger than the reported value of pristine COF-102 ( Wm-1K-1 Liu et al. (2016)). According to the literature, low thermal conductivity of COF-102 primarily originates from vibrational mismatches between B and O atoms Liu et al. (2016). In contrast, such mismatches can be avoided in our derivatives due to the presence of all-carbon linkers.
Following Refs. Ma and Tian (2019); Ma et al. (2021); Hu et al. (2024), to identify origin of the observed behavior, we computed the probability density estimate (PDE) projected onto the plane. PDE is a measure of spatial probability of finding atoms at given spatial locations by projecting all atomic coordinates onto a chosen plane. As shown in Fig. 2(d), COF-102-1 exhibits uniform pore geometries with a high PDE (yellow). In contrast, Figs. 2(e) and 2(f) show that the PDE progressively decreases (shifting toward blue) from COF-102-2 to COF-102-3, indicating increased pore irregularity compared to COF-102-1. Prior studies have established that lower PDE values are associated with greater chain rotation, which reduces overlap of periodic atomic positions, and that such rotational disorder introduces additional vibrational scattering, leading to a weaker temperature dependence in thermal conductivity Ma and Tian (2019); Ma et al. (2021). Based on these findings, we suggest that the reduced PDE in low-density COFs could be attributed to increased linker rotations, which contribute to the anomalous temperature dependence observed in these crystalline structures.
To gain insight into microscopic details, we calculated spectral energy density (SED) to extract phonon dispersion and spectral phonon MFPs Thomas et al. (2010). Phonon dispersions over selected frequency ranges are shown in Fig. 3, while full dispersions are given in Supporting Information Sec. SIV. As shown in Fig. 3(a), COF-102-1 exhibits well-defined phonon modes up to 5 THz, indicating the presence of long-lived thermal phonons. Calculated group velocities near the point of the longitudinal acoustic (LA) and transverse acoustic (TA) branches are 10 km s-1 and 6.5 km s-1, respectively. In contrast, Fig. 3(b) and 3(c) show that COF-102-2 and COF-102-3 exhibit significantly broadened phonon modes in the sub-THz region. Moreover, these structures show reduced group velocities, as shown by the softening of low-frequency acoustic phonon modes compared to COF-102-1. The observed phonon softening and band broadening in them are attributed to previously discussed increased chain rotations, as reported in prior studies Ma and Tian (2019); Ma et al. (2018). We additionally note that, particularly in COF-102-2 and COF-102-3, bor topologies generally exhibit reduced LA group velocities in the low-frequency limit by 30% compared to ctn topologies, along with smaller vibrational density of states (VDOS) peak intensity (see Supporting Information Sec. SV). We attribute our observations to bending deformations of trigonal linkers, which introduce additional vibrational scattering. Previous studies have reported that flexible C–C bonds in carbon-based structures can induce structural deformations, enhancing phonon scattering and thereby reducing phonon lifetimes and thermal conductivity, which is consistent with our observations Zhang (2018a, b); Ying et al. (2021). It is also worth noting that speeds of sound in COF-102 derivatives depend on linker length, in contrast to MOFs. This behavior can be understood by applying a simple 1D chain model as discussed in Ref. Kamencek et al. (2019) (see Supporting Information Sec. SVI for detailed discussions).
[…]
Next, we computed the spectral MFPs following the procedure in Ref. Robbins and Minnich (2015). Figure 3(d) shows the MFP as a function of frequency for all COF-102 derivatives, computed using lifetimes and velocities extracted from the SED results. We observe that COF-102-1 can support sub-THz vibrational modes with MFPs exceeding 1 m, resulting from their high group velocities and long lifetimes. Noting that the specific heat of COF-102-1 is larger than the others owing to its substantially large mass density, such long MFPs, along with large group velocities, imply that these modes may significantly contribute to the observed orders of magnitude increase in thermal conductivity in COF-102-1, as also reported in -conjugated 2D porous polymers Ma et al. (2018).
We further calculated the cumulative thermal conductivity using the non-equilibrium molecular dynamics (NEMD) method. Unlike the EMD method, NEMD predictions are strongly affected by characteristic length of simulated system, known as size effect. Figure 4(a) illustrates the temperature profile of COF-102-1 (ctn), along with a schematic of our simulation setup. In our NEMD simulations, briefly, two fixed regions were placed at the ends of simulation cell, and two Langevin thermostats with a 60 K temperature difference were used to generate heat flux. Finally, thermal conductivity was extracted from the best linear fit of steady-state temperature profile (see Sec. III Methods for details).
Figure 4(b) shows computed thermal conductivity versus , together with cumulative thermal conductivity versus MFPs. According to NEMD simulations, the thermal conductivity of COF-102-1 (ctn) increases from 8.8 Wm-1K-1 to 18.5 Wm-1K-1 as extends from nm to nm. To estimate bulk thermal conductivity from NEMD (), we extrapolated the versus the data, yielding a value of Wm-1K-1, which is in a quantitative agreement with our EMD results (see Supporting Information Sec. SVII).
We next seek to extract thermal conductivity accumulation versus MFPs using the NEMD results by solving an inverse problem described in Ref. Wei et al. (2014). Since the solution of an inverse problem is not unique, we reconstructed all possible solutions using the approach established in prior studies Robbins et al. (2019); Kim et al. (2021). Briefly, this approach enables reconstruction of MFP distribution that satisfies smoothness criteria of priors, while reproducing simulated data by incorporating uncertainty (see Supporting Information Sec. SVIII for descriptions). We obtained the probability distribution of the thermal conductivity accumulation function along with the 95% credible interval using Bayesian inference with a Metropolis–Hastings Markov chain Monte Carlo method Robbins et al. (2019); Kim et al. (2021).
III Methods
@@
We first constructed atomic structures of COF-102 derivatives and optimized their energies using the polymer consistent force field (PCFF) Sun et al. (1994) in Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Plimpton (1995). The detailed structural and physical properties of COF-102 derivatives are provided in Supporting Information Tab. S1 in Sec. II. Further MD calculations with periodic boundary conditions (PBC) applied along all spatial directions were conducted using PCFF potential in LAMMPS.
@@
For EMD simulations, the thermal conductivity is computed using the Green-Kubo formalism Chen (2005).
Here, is system volume, is the Boltzmann constant, is temperature, and is heat flux. The term represents ensemble average of heat current autocorrelation function (HCACF) that decays with time. Thermal conductivity was calculated by integrating HCACF over a correlation time that is sufficiently longer than time at which HCACF decays to zero (see Supporting Information Sec. I for details).
Simulated systems were first relaxed in the NPT, NVT, then NVE ensemble at 300 K for 1 ns. Then, heat flux data were collected in the additional NVE ensemble for 5 ns. To reduce statistical errors, thermal conductivities were averaged over eight independent simulation runs that started with different initial velocities. In all EMD simulations, a time step of 1 fs was used.
@@
In NEMD simulations, we varied domain size along the axial direction () from 33.1 to 198.3 nm, while keeping lateral dimensions equivalent to those used in EMD simulations. The thermal conductivity was computed by using Fourier’s law:
@@
where is heat current, is cross sectional area, and is temperature gradient. System was first relaxed in the NPT ensemble at 300 K for 1 ns. Subsequently, two ends with a length of 2.2 nm were set as fixed regions. Langevin thermostats were applied to heat source and sink regions (length: 8.8 nm) to generate steady-state heat-flux along the axial direction under a temperature gradient for 4 ns; the data (temperature gradient and heat-flux) obtained after initial 1 ns were used to compute the thermal conductivity. Temperatures of heat source and sink regions placed next to each fixed region were set to be at 330 K and 270 K, respectively. A time step of 0.25 fs was used in our entire NEMD calculations.
@@ Acknowledgments
Author contributions
T.K. conceived and supervised the project. S.K. performed the MD calculation and analyzed the data. S.K. and T.K. wrote the manuscript.
Data availability
The authors declare that the data supporting the findings of this study are available within the paper and its Supporting Information files.
Competing interests.
The authors declare no competing interests.

### v2

@@ I Introduction
Covalent organic frameworks (COFs), a class of crystalline porous organic materials, hold promise for various applications such as gas storage and separation Furukawa and Yaghi (2009); Fan et al. (2018), catalysis Wang et al. (2018), optoelectronics Keller and Bein (2021) and energy storage Zhu et al. (2021), owing to their large surface areas, tunable pore sizes, and high thermal stabilities Côté et al. (2005); El-Kaderi et al. (2007); Diercks and Yaghi (2017). Prior extensive studies have suggested that the diverse nature of structure, composition, and functionalities facilitates the tunability of their physical properties for tailored applications Huang et al. (2016); Liu et al. (2021a). For instance, conjugated COFs have shown unique optoelectronic phenomena such as exceptionally high charge carrier mobility Guo et al. (2013) and tunable fluorescence emission as a result of cation addition Ding et al. (2016). Additionally, COFs could exhibit other interesting properties such as facile ion transport Wang et al. (2017); Li et al. (2021) and thermoelectricity Wu et al. (2023); Zhou et al. (2023), which are crucial for energy storage and conversion applications Hu et al. (2020). However, efficient heat dissipation is needed in such applications to ensure device stability Hu and Luo (2019), making it essential to understand heat transport properties in COFs Shakouri (2006); Pop (2010); Babaei et al. (2020).
@@
II Methods
@@
We first constructed atomic structures of COF-102 derivatives and optimized their energies using the polymer consistent force field (PCFF) Sun et al. (1994) in Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Plimpton (1995). The detailed structural and physical properties of COF-102 derivatives are provided in Supporting Information Sec. SI. Further MD calculations with periodic boundary conditions (PBC) applied along all spatial directions were conducted using PCFF potential in LAMMPS.
@@
For EMD simulations, the thermal conductivity was computed using the Green-Kubo formalism Chen (2005).
Here, is system volume, is the Boltzmann constant, is temperature, and is heat current. The term represents ensemble average of heat current autocorrelation function (HCACF) that decays with time. Thermal conductivity was calculated by integrating HCACF over a correlation time that is sufficiently longer than time at which HCACF decays to zero (see Supporting Information Sec. SII for details).
Simulated systems were first relaxed in the NPT, NVT, then NVE ensemble at 300 K for 1 ns. Then, heat current data were collected in the additional NVE ensemble for 5 ns. To reduce statistical errors, thermal conductivities were averaged over eight independent simulation runs that started with different initial velocities. In all EMD simulations, a time step of 1 fs was used. The sizes of simulated structures are summarized in Tab. 1.
@@
In NEMD simulations, we varied domain size along the axial direction () from 33.1 to 198.3 nm, while keeping lateral dimensions equivalent to those used in EMD simulations. The thermal conductivity was computed by using Fourier’s law Liu et al. (2021b); Chakraborty et al. (2020):
@@
where is heat flux and is temperature gradient. System was first relaxed in the NPT ensemble at 300 K for 1 ns. Subsequently, two ends with a length of 2.2 nm were set as fixed regions. Langevin thermostats were applied to heat source and sink regions (length: 8.8 nm) to generate steady-state heat flux along the axial direction under a temperature gradient for 4 ns; the data (temperature gradient and heat flux) obtained after initial 1 ns were used to compute the thermal conductivity. Temperatures of heat source and sink regions placed next to each fixed region were set to be at 330 K and 270 K, respectively. A time step of 0.25 fs was used in our entire NEMD calculations.
III Results and discussion
We studied the thermal transport of COF-102 derivatives using MD simulations. Detailed descriptions of the MD simulations implemented in this work are given in Supporting Information Sec. SII and II Methods. As illustrated in Fig. 1(a), the COF-102 derivatives studied in this work were constructed from two cubic topologies, carbon nitride (ctn) and boracite (bor). They are formed by linking tetrahedral carbon nodes with 3-connected linkers; these nodes are connected to three distinct linkers via C-C bonds. The size of linker depends on the number of phenyl rings it contains. Although both ctn and bor topologies share the same stoichiometry, bor topologies are generally % less dense than ctn topologies (see Supporting Information Sec. SI for details).
Figure 1(b) shows room temperature thermal conductivities of COF-102 derivatives, calculated using the Green-Kubo formalism (Eq. 1). As expected, a positive correlation between thermal conductivity and mass density is clearly visible. We note that the opposite trend is observed between thermal conductivity and pore size (see Supporting Information Fig. S3 in Sec. SII). COF-102-1, in particular, exhibits the highest thermal conductivity ( Wm-1K-1; Wm-1K-1), while the others show thermal conductivities reduced by nearly two orders of magnitude. In all cases, bor topologies yield % lower thermal conductivities than ctn topologies. We here note that COF-102-2 (ctn) is structurally similar to the original COF-102, except that its linkers are formed by substituting all atoms in the boroxine (B3O3) rings with carbon atoms. This leads to enhanced thermal conductivity of Wm-1K-1, which is % larger than the reported value of pristine COF-102 ( Wm-1K-1 Liu et al. (2016)). According to the literature, low thermal conductivity of COF-102 primarily originates from vibrational mismatches between B and O atoms Liu et al. (2016). In contrast, such mismatches can be avoided in our derivatives due to the presence of all-carbon linkers.
Following Refs. Ma and Tian (2019); Ma et al. (2021); Hu et al. (2024), to identify origin of the observed behavior, we computed the probability density estimate (PDE) projected onto the plane. PDE is a measure of spatial probability of finding atoms at given spatial locations by projecting all atomic coordinates onto a chosen plane. As shown in Fig. 2(d), COF-102-1 exhibits uniform pore geometries with a high PDE (yellow). In contrast, Figs. 2(e) and 2(f) show that the PDE progressively decreases (shifting toward blue) from COF-102-2 to COF-102-3, indicating increased pore irregularity compared to COF-102-1. We note that PDE maps for bor topologies exhibit similar trends across the derivatives, as compared to ctn topologies (See Supporting Information Sec. SIV for additional PDE maps of bor structures). Prior studies have established that lower PDE values are associated with greater chain rotation, which reduces overlap of periodic atomic positions, and that such rotational disorder introduces additional vibrational scattering, leading to a weaker temperature dependence in thermal conductivity Ma and Tian (2019); Ma et al. (2021). Based on these findings, we suggest that the reduced PDE in low-density COFs could be attributed to increased linker rotations, which contribute to the anomalous temperature dependence observed in these crystalline structures.
To gain insight into microscopic details, we calculated spectral energy density (SED) using Eq. 2 to extract phonon dispersion and spectral phonon MFPs Thomas et al. (2010). Phonon dispersions over selected frequency ranges are shown in Fig. 3, while full dispersions are given in Supporting Information Sec. SV. As shown in Fig. 3(a), COF-102-1 exhibits well-defined phonon modes up to 5 THz, indicating the presence of long-lived thermal phonons. Calculated group velocities near the point of the longitudinal acoustic (LA) and transverse acoustic (TA) branches are 10 km s-1 and 6.5 km s-1, respectively. In contrast, Fig. 3(b) and 3(c) show that COF-102-2 and COF-102-3 exhibit significantly broadened phonon modes in the sub-THz region. Moreover, these structures show reduced group velocities, as shown by the softening of low-frequency acoustic phonon modes compared to COF-102-1. The observed phonon softening and band broadening in them are attributed to previously discussed increased chain rotations, as reported in prior studies Ma and Tian (2019); Ma et al. (2018). We additionally note that, particularly in COF-102-2 and COF-102-3, bor topologies generally exhibit reduced LA group velocities in the low-frequency limit by 30% compared to ctn topologies, along with smaller vibrational density of states (VDOS) peak intensity (see Supporting Information Sec. SVI). We attribute our observations to bending deformations of trigonal linkers, which introduce additional vibrational scattering. Previous studies have reported that flexible C–C bonds in carbon-based structures can induce structural deformations, enhancing phonon scattering and thereby reducing phonon lifetimes and thermal conductivity, which is consistent with our observations Zhang (2018a, b); Ying et al. (2021). It is also worth noting that speeds of sound in COF-102 derivatives depend on linker length, in contrast to MOFs. This behavior can be understood by applying a simple 1D chain model as discussed in Ref. Kamencek et al. (2019) (see Supporting Information Sec. SVII for detailed discussions).
Next, we computed the spectral MFPs following the procedure in Ref. Robbins and Minnich (2015). Figure 3(d) shows the MFP as a function of frequency for all COF-102 derivatives, computed using velocities extracted from the SED results, along with lifetimes yielded by Lorentzian line-shape fit in Eq. 3. We observe that COF-102-1 can support sub-THz vibrational modes with MFPs exceeding 1 m, resulting from their high group velocities and long lifetimes. Noting that the specific heat of COF-102-1 is larger than the others owing to its substantially large mass density, such long MFPs, along with large group velocities, imply that these modes may significantly contribute to the observed orders of magnitude increase in thermal conductivity in COF-102-1, as also reported in -conjugated 2D porous polymers Ma et al. (2018).
We further calculated the cumulative thermal conductivity using the non-equilibrium molecular dynamics (NEMD) method. Unlike the EMD method, NEMD predictions are strongly affected by characteristic length of simulated system, known as size effect. Figure 4(a) illustrates the temperature profile of COF-102-1 (ctn), along with a schematic of our simulation setup. In our NEMD simulations, briefly, two fixed regions were placed at the ends of simulation cell, and two Langevin thermostats with a 60 K temperature difference were used to generate heat flux. Finally, thermal conductivity was extracted from the best linear fit of steady-state temperature profile computed by Eq. 4 (see Sec. II Methods for details).
Figure 4(b) shows computed thermal conductivity versus , together with cumulative thermal conductivity versus MFPs. According to NEMD simulations, the thermal conductivity of COF-102-1 (ctn) increases from 8.8 Wm-1K-1 to 18.5 Wm-1K-1 as extends from nm to nm. To estimate bulk thermal conductivity from NEMD (), we extrapolated the versus the data, yielding a value of Wm-1K-1, which is in a quantitative agreement with our EMD results (see Supporting Information Sec. SVIII).
We next seek to extract thermal conductivity accumulation versus MFPs using the NEMD results by solving an inverse problem described in Ref. Wei et al. (2014). Since the solution of an inverse problem is not unique, we reconstructed all possible solutions using the approach established in prior studies Robbins et al. (2019); Kim et al. (2021). Briefly, this approach enables reconstruction of MFP distribution that satisfies smoothness criteria of priors, while reproducing simulated data by incorporating uncertainty (see Supporting Information Sec. SIX for descriptions). We obtained the probability distribution of the thermal conductivity accumulation function along with the 95% credible interval using Bayesian inference with a Metropolis–Hastings Markov chain Monte Carlo method Robbins et al. (2019); Kim et al. (2021).
IV Conclusions
Supporting Information
Structural and physical properties of COF-102 derivatives, details of EMD thermal conductivity calculations, XRD patterns, additional PDE maps for bor structures, full phonon dispersion relations using SED analysis, effect of bending deformation on VDOS, dependence of speeds of sound, extrapolation of bulk thermal conductivity in NEMD simulations, MFP reconstruction using Bayesian inference.